assuming that the boric acid, h3po3 present at the end of the borax tritation has a concentration of 0.1 M determine the ph at the end point use this to select an indicator for the titration
Answer #1

Boric acid is H3BO3 and its Ka value (Acid dissociation constant) is 5.8 * 10-10

Boric acid dissociates very weakly and the concentration of H+ ions produced from boric acid is:

[H+] = ( Ka * [H3BO3])1/2 = ( 5.8 * 10-10 * 0.1)1/2 M = 7.62 * 10-6 M

pH = - log[H+] = - log(7.62 * 10-6 ) = 5.12

The best indicator at this pH is bromocresol green.